OpenStudy (anonymous):

How does (2CosX-1)(CosX-3)=0

OpenStudy (anonymous):

This question would be better phrased as when does (2CosX-1)(CosX-3)=0. Regardless you can multiply out to get 2cos^2(x)-6cosx-cosx+3=0 => 2cos^2(x)-7cosx=-3. from there, you can either use acos or use unit circle. knowing cos(pi/3)=1/2, 2(1/4)-7/2=-6/2 or -3. so pi/3 is what your answer is.